Cassidy is 3 times as old as drew. In 8 years Cassidy will only be twice as old is Drew. How old is Cassidy right now?
I know
3d=c
d+8= c/2
and Cassidy is 24 Drew is 8 but I don't know the equation I would use to show my work!
Thanks!
Amie

Answer by ewatrrr(24785)


Hi,
Cassidy is 3 times as old as drew
Let x and 3x represent Drew's and Cassidy's age respectively NOW.
Write as we read***
(3x + 8) = 2(x + 8)
Solving for x
3x + 8 = 2x + 16
x = 8yr , Drew's age NOW, Cassidy is now 24yrs old (3*8)
CHECKING our Answer***
32 = 2*16